OpenStudy (anonymous):

The width of a rectangle is 6 inches less than twice its length. Find the dimension of the rectangle if its area is 108 square inches.

OpenStudy (anonymous):

I just answered this lol. x(2x - 6) = 108 2x² - 6x = 108 x² - 3/2x = 54 + (- 3/2)² x² - 3/2x = 216/4 + 9/4 (x - 3/2)² = 225/4 x - 3/2 = 15/2 x = 18/2 or 9 = 2(9) - 6 = 18 - 6 = 12 9 inches & 12 inches are the dimensions.